OpenStudy (anonymous):

Please help I don't know why I get the wrong volume. The surface area of a ping pong ball is 7.04 in^2. the surface area of a Baseball is 23.75^2. Find the ratio of the volumes of the two.

OpenStudy (anonymous):

My volume is smaller than the radius....

OpenStudy (anonymous):

I will give a medal for best answer

OpenStudy (bibby):

\(SA=4 \pi r^2\\7.04=4 \pi r^2\\0.560=r^2\\r=0.7486\) (ping pong) and such the volume 4/3 pi r^3 = 1.757 do the same thing for the other ball

OpenStudy (anonymous):

so the volume is 1.757? and the volume can be less than the SA?

OpenStudy (bibby):

yeah, I'm trying to logically think about it so I can explain it but I'm too distracted right now

OpenStudy (bibby):

ok, so basically you're comparing apples and oranges surface area is in^2 or whatever volume is in^3

OpenStudy (bibby):

they're entirely different units

OpenStudy (anonymous):

alright so the baseball would be 10.8835?

OpenStudy (bibby):